OpenStudy (anonymous):

Does anyone know how to solve this quad equa someone pls help 4/x+8/x+4=2

OpenStudy (anonymous):

I not understan ,can you rewrite?

OpenStudy (anonymous):

Ok I will try to write it where its legible if its possible :)

OpenStudy (anonymous):

impossible! not a quadriatic equation.

OpenStudy (anonymous):

(4/x)+(8/x+4)= 2 (4(x+4) + 8x)/x(x+4)= 2 4x+16+8x / x^2 +4x =2 12x+16 = 2(x^2 +4x) 12x+16=2x^2+4x 0= 2x^2+4x-12x-16 0= 2x^2-8x-16

OpenStudy (anonymous):

oh ok well its one of my math problems ima see if i can copy and paste it

OpenStudy (anonymous):

if what suzi writes is correct the answer is 4 or -2

OpenStudy (anonymous):

__4___+_8___ =2 lol sorry i dnt know how to use this webiste im new to it but i tried to make it easier idrk if this helps though x x+4

OpenStudy (anonymous):

the quadratic equation is 2x^2-8x-16, u can continue find the x

OpenStudy (anonymous):

like this (4/x)+(8/x+4)=2?

OpenStudy (anonymous):

yes thats it nancy sorry lol

OpenStudy (anonymous):

denominator is: x(x+4) 4(x+4)+8(x)=x(x+4)(2) 4x+16+8x = 2x^2+8x 2x^2+8x-12x-16=0 2x^2-4x-16=0 factor: (2x-8)(x+2) x=-2 or x=4 good luck your home work

OpenStudy (anonymous):

Thank u:)

OpenStudy (anonymous):

I know everyone is right but for some odd reason it says to use the quadratic formula im assumn i have no way of showing yall the way it looks but thanks everyone i really appreciate it goodnight:)

OpenStudy (anonymous):

fans me